# Configurations, Tessellations and Tone Networks

### v4

@@ IV Cyclic Structure as a Basis for Musical Analysis
The relation of the octatonic scales with the tonnetz is as follows (Douthett & Steinbach 1998). Four major chords and four minor chords can be constructed from the constituents of an octatonic scale. For example, from $O_{12}$ one constructs $CM$, AmAm, AMAM, $F_{\sharp}m$, $G_{\flat}M$, $E_{\flat}m$, $E_{\flat}M$, and $Cm$. But these are the chords of the 4pp-octacycle starting at $CM$ in the tonnetz (see the lower left quadrant of Figure 3). One can check that the other two 4pp-octacycles of the tonnetz can be constructed similarly from the other two octatonic scales. The 4pp-octacycles are intimately tied to the three diminished seventh chords, in the same spirit that the four $3p$-hexacycles are tied to the four augmented triads.
By taking the six possible unions of pairs of the four augmented triads we obtain six distinct hexachords. From four of these, one can construct the four $3p$-hexatonic sequences. The remaining two anomalous hexachords are also of interest – they contain the two standard whole-tone scales, each of which contains six notes. These are
$W_{1}=[C,D,E,G_{\flat},A_{\flat},B_{\flat}],\hskip 5.69046ptW_{2}=[D_{\flat},E_{\flat},F,G,A,B].$ (12)
These arrays of six equally spaced whole notes have been exploited by Bartok, Berg, Borodin, Debussy, Janacek, Liszt, Puccini, Rimsky-Korsakov and many other composers.
@@
Numerous examples of the use of melodic and harmonic structures based on cycles of the tonnetz in music of the eighteenth, nineteenth and early twentieth centuries have been noted (Cohn 2012 and references cited therein) in works of composers such as Mozart, Beethoven, Schubert, Schumann, Chopin, Wagner, Liszt, Bruckner, Brahms, Mahler and Strauss.   [^12] Such use is not obvious to the classically trained ear since the progressions associated with tonnetz cycles do not form part of the training in basic harmony taught in everyday music theory. Nor would the composers listed above have been conscious of their own use of the structural determinants implicit in the cycles of the tonnetz. If the unconscious mind will tap the resources of the mathematical structures embedded in the geometry of the tonnetz then surely it will seek to avail itself of the full richness of these resources.
One advantage of the use of the twelve 2pp-hexacycles, the four 3pp-hexacycles, the twelve $1p$-octacycles, the three $4p$-octacycles, and so on, all together, in their manifestation as elements of a single cycle space, is that collectively the cycles of the tonnetz incorporate a diversity of modality sufficient to enable them to act as a basis both for the composition of music and for its analysis; thus, in our view the general scheme of pan-triadic musicology should involve all the cycles of the tonnetz on essentially an equal footing, the shorter ones being given preference only to the extent that (a) they are more likely to emerge as being relevant in the context of shorter passages of music and (b) the longer ones can be assembled by a process of accretion (i.e., surgical insertion) from shorter ones.
@@
As an example, let us return to the problem of understanding the resolution of the Tristan chord. Wagner offers a well known ‘tentative’ resolution of his permuted $G_{\sharp}m^{\,6}$ into an $E^{\,7}$, shown in Figure 5. Chopin, on the other hand, resolves the chord into a $B_{\flat}^{\,7}$, as we see in Figure 6. Chopin’s resolution is definitive: it brings that lengthy first section of the Ballade to a decisive close, and paves the way forward at the più animato with a change of tempo and the introduction of new material and a new atmosphere. Wagner’s resolution is inconclusive, in line with the ever shifting sensibilities of the protagonists.
@@
A glance at Figure 2 is revealing. Chopin’s resolution takes the initial $G_{\sharp}m$ along the path of a $2p$-decacycle to its polar opposite $B_{\flat}M$. From $G_{\sharp}m$ we follow the ‘floppy bow tie’ trajectory to $BM$, then to $E_{\flat}m$, then across to $E_{\flat}M$, up to $Gm$, and finally to $B_{\flat}M$ at the opposite side of the decagon. Arrival at the polar opposite gives this resolution its firmness. The situation with Wagner’s ambivalent resolution is less decisive: he merely moves the chord to an adjacent one, one step clockwise along the unique 3pp-hexacycle to which both chords belong from $G_{\sharp}m$ to $EM$. The simplicity of this transformation suggests why its character is so unsettled: nothing has changed. One is moving sideways, from indecisiveness to indecisiveness.
Indeed, we submit that by studying the relative positions of the two underlying triads on the tonnetz and how the one chord can be reached from the other along a trajectory of a cycle one is offered insight into the musicality of these transitions.
@@
There is a substantial literature addressing the issue of how four-note chords might be accommodated into tonnetz-like schemes (Childs 1998, Douthett & Steinbach 1998, Cohn 2012, Tymoczko 2011, 2012, Nuño 2021). Such schemes have variously included dominant sevenths, minor sixths and minor sevenths, as well as other chords such as diminished sevenths and French sixths.  [^13] Some tetrachords are ambiguous – for example, $Cm^{7}$ and $E_{\flat}M^{6}$ are the same chord modulo a permutation. What is the underlying triad, $Cm$ or $E_{\flat}M$? The same is true of the minor sixth, which can also function as a half-diminished seventh. The virtue of such chords is that they can function in either capacity: they are ambidextrous, so one needs to take into account the context in which they appear to determine their function. Two historical precedents for this well known idea are Rameau’s notion of double emploi and Gottfried Weber’s Mehrdeutigkeit.  We present our own approach to tetrachord tonnetze later in Section X. Nonetheless, tetrachord relations can be given an illuminating partial analysis in terms of the triads underpinning them.  [^14] We draw attention to Cohn (2012) in Chapter 7, ‘Dissonance’, at the section ‘Reduction to a Triadic Subset’ (pp. 142-145); NB Cohn’s remark at p. 144, ‘Wagner often treats minor triads and their $\phi^{7}$ supersets as interchangeable in his late music.’
@@
Figure 6: Chopin, Ballade in $G$-minor, Opus 23, excerpt. The Tristan chord appears prominently in the left hand, as an arpeggio, fortissimo, at Wagner’s pitches, but in the enharmonically equivalent form $E_{\sharp},B,D_{\sharp},G_{\sharp}]$, a permutation of the $G_{\sharp}m^{\,6}$ chord. The descending figure in the right hand touches the same notes. The resolution is into a $B_{\flat}^{\,7}$.
@@
A prominent example of a resolution of a Tristan-genus chord is found in the Finale ‘Adagio Lamentoso’ of Tchaikovsky’s Symphony No. 6 in B-Minor, Opus 74, ‘Pathetique’. Tchaikovsky had been studying Wagner’s works and evidently wished to make use of the novel harmonic structures implicit in this material. Tchaikovsky’s modified version of the Tristan chord, heard memorably in the strings at the outset of the Adagio Lamentoso carries the full force of Wagner’s permutation. Recall that Wagner’s chord is $F,B,E_{\flat},A_{\flat}]$, whereas Tchaikovsky, in line with the overall key of his symphony uses a $Bm^{6}$ chord, in the permutation $D,G_{\sharp},B,F_{\sharp}]$, as in Figure 7, resolving into an $F_{\sharp}^{\,7}$. If we transpose this passage down three semitones to facilitate comparison with Wagner and Chopin the result is a modified Tristan chord in the form $B,F,A_{\flat},E_{\flat}]$, which resolves to $E_{\flat}^{\,7}$.
@@
However, Tchaikovsky’s resolution to $E_{\flat}M$, since it is only across a $2p$-hexacycle rather than a $2p$-decacycle, as in the case of Chopin, thus has a more perfunctory character, that of a plagal cadence, and hence its desolation. It is interesting to note, in contrast, that in the $G$-minor Ballade much of the passage leading up to the Tristan chord is based, in part, on the key of $E_{\flat}M$ (see the analysis of Chopin’s Opus 23 undertaken in Cohn 2012 at pp. 96-98).
@@
One can ask, did Wagner himself make use of Tchaikovsky’s version of the Tristan chord? Was Tchaikovsky aware of this? The reply is ‘yes’ and ‘very likely’. Tchaikovsky’s chord can be found in Parsifal, early in Act 3, at the passage where the horns mysteriously announce the return of Parsifal to the now languishing Kingdom of the Holy Grail, dressed in black armour and carrying the Sacred Spear.
The Parsifal motiv, which in the first act took the form of a simple fanfare involving a triplet figure leading from a $B_{\flat}M$ to an $E_{\flat}M$ has evolved (Figure 8) into a fanfare on $B_{\flat}m^{6}$ leading to a devastating $D_{\flat}m^{6}$. It is this $D_{\flat}m^{6}$, which Wagner presents in the form $E_{\natural},B_{\flat},D_{\flat},A_{\flat}]$, that is isomorphic to the $Bm^{6}$ that Tchaikovsky presents in the form $D,G_{\sharp},B,F_{\sharp}]$ in the Adagio Lamentoso.
Wagner’s version of the chord is pitched a tone higher than that of Tchaikovsky, but lowered by an octave with the $E_{\natural}$ doubled in the bass to add darkness and weight. It does not resolve to a dominant seventh, but rather, and only tentatively, to a $G_{\flat}m^{6}$ in the form $E_{\flat},B_{\flat\flat},D_{\flat},G_{\flat}]$, which is a Tristan chord. Despite the surrounding sense of gloom and oppression, Wagner’s resolution is ultimately conventional – in keeping with Parsifal’s character – i.e., up a fourth from $D_{\flat}m^{6}$, in line with the basic Parsifal motiv – and hence consoling. The Tristan chord then resolves further, with more certainty, to another Tristan chord a tone lower given by an $Em^{6}$ in the form $C_{\sharp},G,B,E]$.
Figure 8: Wagner, Parsifal, excerpt from Act 3. The Parsifal leitmotiv, which at Parsifal’s entry in Act 1 is a simple fanfare leading from $B_{\flat}M$ to $E_{\flat}M$, has evolved at Parsifal’s Act 3 entry into $B_{\flat}m^{6}$ leading to $D_{\flat}m^{6}$, which Wagner presents in the permutation $\{E_{\natural},B_{\flat},D_{\flat},A_{\flat}\}$, isomorphic to Tchaikovsky’s $Bm^{6}$ in the form $\{D,G_{\sharp},B,F_{\sharp}\}$ appearing at the outset of the Adagio Lamentoso.
@@
The diminished seventh $G,E,A_{\sharp},C_{\sharp}]$ that finishes off this sequence of four minor sixths representing Gurnemanz’s take on the unexpected guest is merely one semitone away from a $Gm^{6}$, and hence for Wagner’s dramatic purposes serves well as a point of exit from the decacycle. This view is not inconsistent with the idea that the Hamiltonian cycle circumnavigating the tonnetz is also playing a role – for not only is $Em$ opposite to $B_{\flat}m$ in the Hamiltonian cycle, one also finds that $Gm$ is opposite to $D_{\flat}m$. Hence the chords $B_{\flat}m$, $D_{\flat}m$, $Em$, GmGm are equally spaced around the tonnetz and the names of these chords spell out the notes of the diminished seventh on which the sequence ends.
Now, as for whether Tchaikovsky knew about Wagner’s use of the Adagio Lamentoso chord – well, we know that Tchaikovsky possessed a piano reduction of Parsifal and had played through it a number of times – so he would have been well aware of Wagner’s use of the Tristan chord in that opera.  [^15] One can only speculate about the nature of Tchaikovsky’s thought processes as he composed his final symphony. Tchaikovsky’s chord can also be found at the strident opening of Chopin’s Scherzo in $B$-Minor, Opus 20, in the form $G,C_{\sharp},E,B]$, resolving to $A_{\sharp},C_{\sharp},E,F_{\sharp}]$, both with some doubling. Thus, we have $Em^{6}$ moving to $F_{\sharp}^{\,7}$, the first chord being in Tchaikovsky’s permutation and the second being, unusually, in the first inversion. Putting this into Wagner’s Tristan pitch levels, we have $G_{\sharp}m^{6}$ moving to $A_{\sharp}^{\,7}$, which we recognize as enharmonically equivalent to the resolution in the $G$-minor Ballade at the piú animato; though here since the dominant seventh appears in the first inversion it has a different character, demanding immediate further resolution to the tonic $D_{\sharp}m$, which in the key of the Scherzo is $Bm$.
@@ V Pentatonic Tone Network
‘Musical keys and modulations do appear to arouse certain respondent moods or emotions. But is this, primarily, a matter of historical convention, of schooled expectation? What makes a minor third ‘sad’? Is G-minor, in the Western scale, intrinsically triste (and just what could such a statement signify?), or does its desolation stem from the use Mozart made of it in his great Quintet K 516? What of the reflexes of sensibility motivated by keys, pitch, chordal blocks in non-Western tonic systems? Is a pentatonic structure any less universal than ours?’ —George Steiner (1997), Errata
According to Piston (1985), ‘The pentatonic scale, whose intervallic pattern corresponds to that of the black keys on the piano, is very ancient, having been used in the music of Oriental cultures perhaps longer even than the diatonic scale in the West. It is also the scale used in many European folksongs, particularly those of the British Isles.’ Examples of the use of pentatonic scales can also be found in the works of many of the well-known modern Western composers, including Bartok, Chopin, Debussy, Mahler, Ravel and Stravinsky, to name a few.  [^16] Piston/DeVoto’s characterization of the broad geographical range of the pentatonic scale may or may not be strongly research-backed. It would be interesting to know to what extent these remarks can be substantiated by recent ethnomusicological literature. Could there be an analogue of the tonnetz based on the pentatonic scale? Although some bare rudiments of the system of major and minor chords can be glimpsed in the pentatonic system, nothing like the rich dodecaphonic system of Western music exists, so it is not easy to imagine off hand what form a pentatonic tonnetz might take.
@@
Instead of triads, the major and minor chords are clusters of two and three notes, respectively. In pentatonic music these clusters possess a pleasing demeanor – none sound dissonant or distressing. If we take the pentatonic scale to be the notes C,D,E,G,AC,\,D,\,E,\,G,\,A, the ten two-note clusters are $\{CD\},\,\{CE\},\,\{CG\},\,\{CA\},\,\{DE\},\,\{DG\},\,\{DA\},\,\{EG\},\,\{EA\}$, $\{GA\}$, and the ten three-note clusters are $\{ACD\}$, $\{CDE\}$, $\{DEG\}$, $\{EGA\}$, $\{GAC\}$, $\{CDG\}$, $\{DEA\}$, $\{EGC\}$, $\{GAD\}$, $\{ACE\}$.
@@
Another example can be found in the Molto Adagio at the conclusion of the fugue in Bach’s Toccata and Fugue in $D$ Minor, where a permutation of the cluster $B_{\flat}CD]$ is accompanied by an $F$, leading to the cathartic minor plagal cadence at the terminus. But the occurrence of such tone clusters tends to be sporadic in the extended common practice and mostly for effect rather syntax. Can one make music from clusters alone?
@@
In fact, the musical labelling of the Desargues configuration shown in Figure 11 is in some respects even more memorable than the traditional geometric labelling of Figure 10. The five tones of the pentatonic scale yield ten two-note ‘major’ clusters and ten three-note ‘minor’ clusters. The pitches of these tones may vary from culture to culture even if the combinatorial geometry of the configuration is invariant. One could even consider a form of pentatonic equal temperament by setting
$C=1,\,\,D=2^{1/5},\,\,E=2^{2/5},\,\,G=2^{3/5},\,\,A=2^{4/5}.$ (14)
We refer to the resulting clusters of two and three tones as the ‘chords’ of pentatonic music. Each major chord can be found in three minor chords and each minor chord contains three major chords.
For example, the major chord $CD$ can be found in the three minor chords $CDE$, CDGCDG, and $CDA$, whereas the minor chord $CDE$ contains the major chords $CD$, CECE, and $DE$, as one sees in Figure 11. Here, to ease the notation we drop the braces around unordered chords. Thus we obtain a three to three map between the major chord set and the minor chord set. Figure 12 shows on the left how the chords of the pentatonic tone network can be arranged in the form of a Levi graph, of which another view can be found on the right in which some of the other cycles can be easily discerned. What is remarkable is the overall affinity of the underlying combinatorial geometries of the pentatonic tonnetz and the Eulerian tonnetz.

### v5

@@ IV Cyclic Structure as a Basis for Musical Analysis
The relation of the octatonic scales with the tonnetz is as follows (Douthett & Steinbach 1998). Four major chords and four minor chords can be constructed from the constituents of an octatonic scale. For example, from $O_{12}$ one constructs $CM$, AmAm, AMAM, $F_{\sharp}m$, $G_{\flat}M$, $E_{\flat}m$, $E_{\flat}M$, and $Cm$. But these are the chords of the 4pp-octacycle starting at $CM$ in the tonnetz (see lower left quadrant of Figure 3). The other two 4pp-octacycles of the tonnetz can be constructed similarly from the other two octatonic scales. The 4pp-octacycles are tied to the three diminished seventh chords, in the same spirit that the four $3p$-hexacycles are tied to the four augmented triads. By taking the six possible unions of pairs of the four augmented triads we obtain six distinct hexachords. From four of these, one constructs the four $3p$-hexatonic sequences. The remaining two hexachords contain the standard whole-tone scales, each of which contains six notes. These are $W_{1}=[C,D,E,G_{\flat},A_{\flat},B_{\flat},C]$ and $W_{2}=[D_{\flat},E_{\flat},F,G,A,B,D_{\flat}]$. These arrays of six equally spaced whole notes have been exploited by Bartok, Berg, Borodin, Debussy, Janacek, Liszt, Puccini, Rimsky-Korsakov and many other composers.
@@
Numerous examples of the use of melodic and harmonic structures based on cycles of the tonnetz in music of the common practice period have been noted (Cohn 2012 and references cited therein) in works of composers such as Mozart, Beethoven, Schubert, Schumann, Chopin, Wagner, Liszt, Bruckner, Brahms, Mahler and Strauss. Such use is not obvious to the classically trained ear since the progressions associated with tonnetz cycles do not form part of the training in basic harmony taught in everyday music theory. Nor would the composers listed have been conscious of their own use of the structural determinants implicit in the cycles of the tonnetz. If the unconscious mind will tap the resources of the mathematical structures embedded in the geometry of the tonnetz then surely it will seek to avail itself of the full richness of these resources. One advantage of the use of the twelve 2pp-hexacycles, the four 3pp-hexacycles, the twelve $1p$-octacycles, the three $4p$-octacycles, and so on, all together, in their manifestation as elements of a single cycle space, is that collectively the cycles of the tonnetz incorporate a diversity of modality sufficient to enable them to act as a basis both for the composition of music and for its analysis; thus, in our view the general scheme of pan-triadic musicology should involve all the 5409 cycles of the tonnetz. See the Appendix for an enumeration of the different types of cycles of the tonnetz.
@@
As an example of the analysis of cyclic structure, let us return to the problem of understanding the resolution of the Tristan chord. Wagner offers a well known ‘tentative’ resolution of his permuted $G_{\sharp}m^{\,6}$ to $E^{\,7}$, as shown in Figure 5. Chopin, on the other hand, resolves the same chord to $B_{\flat}^{\,7}$, as we see in Figure 6. Chopin’s resolution is definitive: it brings that lengthy first section of the Ballade to a decisive close, and paves the way forward at the più animato with a change of tempo and the introduction of new material and a new atmosphere. Wagner’s resolution is inconclusive, in line with the ever shifting sensibilities of the protagonists.
@@
In this connection a glance at Figure 2 is revealing. One sees that Chopin’s resolution takes the initial $G_{\sharp}m$ along the path of a $2p$-decacycle to its polar opposite $B_{\flat}M$. From $G_{\sharp}m$ we follow the ‘floppy bow tie’ trajectory to $BM$, then to $E_{\flat}m$, then across to $E_{\flat}M$, up to $Gm$, and finally to $B_{\flat}M$ at the opposite side of the decagon. It seems that it is arrival at the polar opposite that gives this resolution its firmness. The situation with Wagner’s ambivalent resolution is less decisive: he merely moves the chord to an adjacent one, one step clockwise along the unique 3pp-hexacycle to which both chords belong from $G_{\sharp}m$ to $EM$. The simplicity of this transformation suggests why its character is so unsettled: nothing has changed. One is moving sideways, from indecisiveness to indecisiveness. Indeed, we submit that by studying the relative positions of the two underlying triads on the tonnetz and how the one chord can be reached from the other along a trajectory of a cycle one is offered insight into the musicality of these transitions.
@@
There is a substantial literature addressing the issue of how four-note chords might be accommodated into tonnetz-like schemes (Childs 1998, Douthett & Steinbach 1998, Cohn 2012, Tymoczko 2011b, 2012, Nuño 2021). Such schemes have variously included dominant sevenths, minor sixths and minor sevenths, as well as other chords such as diminished sevenths and French sixths.  [^18] Note that $Cm^{7}$ and $E_{\flat}M^{6}$ are the same chord modulo a permutation. What is the underlying triad, $Cm$ or $E_{\flat}M$? The same is true of the minor sixth, which can also function as a half-diminished seventh. The virtue of such chords is that they can function in either capacity: they are ambidextrous, so one needs to take into account the context in which they appear to determine their function. Two historical precedents for this well known idea are Rameau’s notion of double emploi and Gottfried Weber’s Mehrdeutigkeit.  We present our own approach to tetrachord tonnetze in Section X. Nonetheless, in the meantime we take advantage of the fact that tetradic relations can be given an illuminating partial analysis in terms of the triads underpinning them.  [^19] We draw attention to Cohn (2012) in Chapter 7, ‘Dissonance’, at the section ‘Reduction to a Triadic Subset’ (pp. 142-145); NB Cohn’s remark at p. 144, ‘Wagner often treats minor triads and their $\phi^{7}$ supersets as interchangeable in his late music.’
@@
Figure 6: Chopin, Ballade in $G$-minor, Opus 23, excerpt. The Tristan chord appears in the left hand, as an arpeggio, fortissimo, at Wagner’s pitches, but in the enharmonically equivalent form $E_{\sharp},B,D_{\sharp},G_{\sharp}]$. The descending figure in the right hand touches the same notes and is mirrored by the violins in Tristan und Isolde at the climax of the prelude. Chopin’s resolution is to $B_{\flat}^{\,7}$.
@@
A prominent example of a resolution of a Tristan-genus chord is found in the Finale ‘Adagio Lamentoso’ of Tchaikovsky’s Symphony No. 6 in B-Minor, Opus 74, ‘Pathetique’. Tchaikovsky had been studying Wagner’s works and evidently wished to make use of the novel harmonic structures implicit in this material. Tchaikovsky’s modified version of the Tristan chord, heard memorably in the strings at the outset of the Adagio Lamentoso carries the full force of Wagner’s permutation. Recall that Wagner’s chord is $F,B,E_{\flat},A_{\flat}]$, whereas Tchaikovsky, in line with the overall key of his symphony uses a $Bm^{6}$ chord, in the permutation $D,G_{\sharp},B,F_{\sharp}]$, as in Figure 7, resolving into $F_{\sharp}^{\,7}$. If we transpose this passage down three semitones to facilitate comparison with Wagner and Chopin the result is a modified Tristan chord in the form $B,F,A_{\flat},E_{\flat}]$, which resolves to $E_{\flat}^{\,7}$.
@@
However, Tchaikovsky’s resolution to $E_{\flat}M$, since it is only across a $2p$-hexacycle rather than a $2p$-decacycle, as in the case of Chopin, thus has a more perfunctory character, that of a plagal cadence, and hence its desolation. We note, in contrast, that in the $G$-minor Ballade much of the passage leading up to the Tristan chord is based, in part, on the key of $E_{\flat}M$ (see the analysis of Chopin’s Opus 23 undertaken in Cohn 2012 at pp. 96-98).
@@
One can ask, did Wagner himself make use of Tchaikovsky’s version of the Tristan chord? Was Tchaikovsky aware of this? The reply is ‘yes’ and ‘very likely’. Tchaikovsky’s chord can be found in Parsifal, early in Act 3, at the passage where the horns mysteriously announce the return of Parsifal to the now languishing Kingdom of the Holy Grail, dressed in black armour and carrying the Sacred Spear. The Parsifal motif, which in the first act took the form of a simple fanfare involving a triplet figure leading from $B_{\flat}M$ to $E_{\flat}M$ has evolved (Figure 8) into a fanfare on $B_{\flat}m^{6}$ leading to a devastating $D_{\flat}m^{6}$. It is this $D_{\flat}m^{6}$, which Wagner presents in the form $E_{\natural},B_{\flat},D_{\flat},A_{\flat}]$, that is isomorphic to the $Bm^{6}$ that Tchaikovsky presents in the form $D,G_{\sharp},B,F_{\sharp}]$ in the Adagio Lamentoso.
Wagner’s version of the chord is pitched a tone higher than that of Tchaikovsky, but lowered by an octave with the $E_{\natural}$ doubled in the bass to add darkness and weight. It does not resolve to a dominant seventh, but rather, and only tentatively, to a $G_{\flat}m^{6}$ in the form $E_{\flat},B_{\flat\flat},D_{\flat},G_{\flat}]$, which is a Tristan chord. Despite the surrounding sense of gloom and oppression, Wagner’s resolution is ultimately conventional – in keeping with Parsifal’s character – i.e., up a fourth from $D_{\flat}m^{6}$, in line with the basic Parsifal motif – and hence consoling. The Tristan chord then resolves further, with more certainty, to another Tristan chord a tone lower given by an $Em^{6}$ in the form $C_{\sharp},G,B,E]$.
Figure 8: Wagner, Parsifal, excerpt from Act 3. The Parsifal leitmotif, which at Parsifal’s entry in Act 1 is a simple fanfare leading from $B_{\flat}M$ to $E_{\flat}M$, has evolved at Parsifal’s Act 3 entry into $B_{\flat}m^{6}$ leading to $D_{\flat}m^{6}$, which Wagner presents in the permutation $\{E_{\natural},B_{\flat},D_{\flat},A_{\flat}\}$, isomorphic to Tchaikovsky’s $Bm^{6}$ in the form $\{D,G_{\sharp},B,F_{\sharp}\}$ appearing at the outset of the Adagio Lamentoso.
@@
The diminished seventh $G,E,A_{\sharp},C_{\sharp}]$ that finishes off this sequence of four minor sixths representing Gurnemanz’s take on the unexpected guest is merely one semitone away from a $Gm^{6}$, and hence for Wagner’s dramatic purposes serves well as a point of exit from the decacycle. This view is not inconsistent with the idea that the Hamiltonian cycle circumnavigating the tonnetz is also playing a role – for not only is $Em$ opposite to $B_{\flat}m$ in the Hamiltonian cycle, one also finds that $Gm$ is opposite to $D_{\flat}m$. Hence the chords $B_{\flat}m$, $D_{\flat}m$, $Em$, GmGm are equally spaced around the perimeter of the tonnetz and the names of these chords spell out the notes of the diminished seventh on which the sequence ends.
Now, as for whether Tchaikovsky knew about Wagner’s use of the Adagio Lamentoso chord – well, we know that Tchaikovsky possessed a piano reduction of Parsifal and had played through it a number of times – so, indeed, he would have been well aware of Wagner’s use of the Tristan chord in that opera. Is it possible that Tchaikovsky got his idea elsewhere? Yes! Tchaikovsky’s chord can also be found at the strident opening of Chopin’s Scherzo in $B$-Minor, Opus 20, in the form $G,C_{\sharp},E,B]$, resolving to $A_{\sharp},C_{\sharp},E,F_{\sharp}]$, both with some doubling. Thus, we have $Em^{6}$ moving to $F_{\sharp}^{\,7}$, the first chord being in Tchaikovsky’s permutation and the second being, unusually, in the first inversion. Putting this into Wagner’s Tristan pitch levels, we have $G_{\sharp}m^{6}$ moving to $A_{\sharp}^{\,7}$, which we recognize as enharmonically equivalent to the resolution in the $G$-minor Ballade at the piú animato; though here since the dominant seventh appears in the first inversion it has a different character, demanding immediate further resolution to the tonic $D_{\sharp}m$, which in the key of the Scherzo is $Bm$.
@@ V Pentatonic Tone Network
‘Musical keys and modulations do appear to arouse certain respondent moods or emotions. But is this, primarily, a matter of historical convention, of schooled expectation? What makes a minor third ‘sad’? Is G-minor, in the Western scale, intrinsically triste (and just what could such a statement signify?), or does its desolation stem from the use Mozart made of it in his great Quintet K 516? What of the reflexes of sensibility motivated by keys, pitch, chordal blocks in non-Western tonic systems? Is a pentatonic structure any less universal than ours?’ —George Steiner, Errata
According to Piston (1985), ‘The pentatonic scale, whose intervallic pattern corresponds to that of the black keys on the piano, is very ancient, having been used in the music of Oriental cultures perhaps longer even than the diatonic scale in the West. It is also the scale used in many European folksongs, particularly those of the British Isles.’ Examples of the use of pentatonic scales can also be found in the works of many of the well-known modern Western composers, including Bartok, Chopin, Debussy, Mahler, Ravel and Stravinsky, to name a few.  [^20] Piston/DeVoto’s characterization of the broad geographical range of the pentatonic scale may or may not be strongly research-backed. It would be interesting to know to what extent the claim there can be substantiated by recent ethnomusicological literature. Could there be an analogue of the tonnetz based on the pentatonic scale? Although some bare rudiments of the system of major and minor chords can be glimpsed in the pentatonic system, nothing like the rich dodecaphonic system of Western music exists, so it is not easy to imagine off hand what form a pentatonic tonnetz might take.
@@
Instead of triads, the major and minor chords are clusters of two and three notes, respectively. In pentatonic music these clusters possess a pleasing demeanor – none sound particularly distressing. If we take the pentatonic scale to be the notes C,D,E,G,AC,\,D,\,E,\,G,\,A, the ten two-note clusters are $\{CD\},\,\{CE\},\,\{CG\},\,\{CA\},\,\{DE\},\,\{DG\},\,\{DA\},\,\{EG\},\,\{EA\}$, $\{GA\}$, and the ten three-note clusters are $\{ACD\}$, $\{CDE\}$, $\{DEG\}$, $\{EGA\}$, $\{GAC\}$, $\{CDG\}$, $\{DEA\}$, $\{EGC\}$, $\{GAD\}$, $\{ACE\}$.
@@
Another example can be found in the Molto Adagio at the conclusion of the fugue in Bach’s Toccata and Fugue in $D$ Minor, where a permutation of the cluster $B_{\flat}CD]$ is accompanied by an $F$, leading to the cathartic minor plagal cadence at the terminus. But the occurrence of such diatonic tone clusters tends to be sporadic in the extended common practice and mostly for colour and effect rather than logic and syntax. Can one make music from clusters alone?
@@
In fact, the musical labelling of the Desargues configuration shown in Figure 11 is in some respects even more memorable than the traditional geometric labelling of Figure 10. The five tones of the pentatonic scale yield ten two-note ‘major’ clusters and ten three-note ‘minor’ clusters. The pitches of these tones may vary from culture to culture even if the combinatorial geometry of the configuration is invariant. One could even consider a form of pentatonic equal temperament by setting $C=1$,  $D=2^{1/5}$,  $E=2^{2/5}$,  $G=2^{3/5}$,  $A=2^{4/5}$. We refer to the resulting clusters of two and three tones as the ‘chords’ of pentatonic music. Each major chord can be found in three minor chords and each minor chord contains three major chords. For example, the major chord $CD$ can be found in the three minor chords $CDE$, CDGCDG, and $CDA$, whereas the minor chord $CDE$ contains the major chords $CD$, CECE, and $DE$, as one sees in Figure 11. Here, to ease the notation we drop the braces around unordered chords. Thus we obtain a three to three map between the major chord set and the minor chord set.
Figure 12 shows on the left how the chords of the pentatonic tone network can be arranged in the form of a Levi graph, of which another view can be found on the right in which some of the other cycles can be easily discerned. What is remarkable is the overall affinity of the underlying combinatorial geometries of the pentatonic tonnetz and the Eulerian tonnetz. The results are summarized in Proposition 3. The pentatonic tonnetz can be readily used as a compositional tool. In Figure 13 we give an example of such a composition.
Looking at the Desargues configuration in Figures 10 and 11 one might be tempted to assign some special significance to the point of perspective $V$ or in the musical picture the chord $GA$. But this is illusory – any of the ten points can be taken as a point of perspective, from which the two associated triangles can be identified and the line of perspective.  [^21] Nonetheless, once the symmetry is broken by the choice of a point of perspective the other points line up to it in a specific way with certain relations among them. Thus, the choice of a preferred chord, such as $GA$, as the point of perspective is analogous to a choice of key. No one choice is preferred a priori, but once a choice has been made then the music associated with that chord involves a specific set of relations between that chord and the other chords. This is similar to the way in which in common practice tonality no specific key is preferred a priori to another – they are all on an equal footing – but once a key has been chosen then much else follows. The two-letter ‘major’ chords determine ten major keys; and since the configuration is self-dual, the ten three-letter ‘minor’ chords determine ten minor keys. A more symmetrical way of drawing the Desargues configuration can be found in figure 2.3a of Coxeter (1974) in which two different points of perspective, and hence two different lines of perspective, are treated in a manifestly symmetric way.